March 4-7

– Four stores provided ~41 hours of luminosity
– Store 8549 quenched due to a kicker problem
– A Switchyard septa tripped off due to a vacuum problem
– The Tevatron quenched during shot setup due to another kicker problem
– Linac and Operator personnel tuned up the I Source
– DZero required eight-hour access for repairs
– Tevatron and Antiproton Source personnel conducted studies
– Due to ComEd work, Operations turned off the accelerator complex for machine upgrades and repairs

* The integrated luminosity for the period from Feb. 21- March 7 was 52.1 inverse picobarns. NuMI was off to replace its target.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts